The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs.
Answer by Paul(988)

Length = 4x-8
Base = x
4x-8+x=29
5x=37
x=7.4
4(7.4)-8=21.6
21.6/2 = 10.8
The length of 1 equal side is 10.8 cm, and the base is 7.4 cm.
Paul.